OpenStudy (anonymous):

@satellite...can you help me please pretty please on these complex polynomials?

OpenStudy (anonymous):

x^3-6x^2+25x

OpenStudy (anonymous):

and x^3+9x^2+22x

OpenStudy (anonymous):

what are you supposed to do?

OpenStudy (anonymous):

factor them out, like we did, but when I took out the x, i found it difficult to factor the inside

OpenStudy (anonymous):

use the quadratic formula?

OpenStudy (anonymous):

okay, so if i get something like (6 plus or minus √-64)/2 how would it look simplified?

OpenStudy (anonymous):

let me check

OpenStudy (anonymous):

\[x^3-6x^2+25x\] \[x(x^2-6x+25)\] we can solve \[x^2-6x+25=0\] to factor easy to complete the square \[x^2-6x-25=0\] \[x^2-6x=-25\] \[(x-3)^2=-25+9=-16\] \[x-3=\pm4i\] \[x=3\pm 4i\]

OpenStudy (anonymous):

so if you want the zeros they are \[\{0,3+4i,3-4i\}\] and if you want to factor it would be \[x(x-(3+4i))(x-(3-4i))\]

OpenStudy (anonymous):

now if you got \[\frac{6\pm\sqrt{-64}}{2}\] that is the same because it is \[\frac{6\pm8i}{2}=3\pm4i\]

OpenStudy (anonymous):

oh yeah ^_^

OpenStudy (anonymous):

x^3+9x^2+22x ill try doing this one, and can u check my answer?

OpenStudy (anonymous):

sure i will do it and wait

OpenStudy (anonymous):

i mean first part is just write as \[x(x^2+9x+22)\] that is easy. then solve the quadratic

OpenStudy (anonymous):

okay after u made it equal to -25 where did the 9 come from?

OpenStudy (anonymous):

completing the square i turned \[x^2-6x\] into \[x^2-6x+9=(x-3)^2\] by adding 9, so i added it on the right as well

OpenStudy (anonymous):

so what would be the final answer? i only have like 7 min left to turn it in :(